# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v2

@@ 5.3.1 Theoretical Contributions
Potential Benefits and Risks. For some individuals, particularly those with strong existing social networks and positive psychological profiles, companion chatbots may serve as a beneficial supplement to human social interaction. These users may experience higher social confidence in parallel with their chatbot use. Our data suggest that neurotic individuals, who tend to be lonely and socially isolated, may experience reduced loneliness when spending more time with chatbots, but previous research suggests that more socially isolated individuals are at greater risk of problematic use of online services [Reer et al.(2019), O’Day and Heimberg(2021)]. As our model should be interpreted cautiously, the risk of chatbots replacing human interactions and potentially exacerbating loneliness and social withdrawal should be considered.
Implications for HCI Design and Practice. Our findings on problematic use suggest the need for built-in mechanisms to detect patterns of potentially harmful usage, provide user feedback about usage patterns, and implement interventions when problematic patterns emerge. Additionally, offering resources for maintaining healthy human relationships alongside chatbot use may prove beneficial if they can help improve the user’s social support.
Further, if the user’s traits can be collected, designing adaptive interfaces for social chatbots based on user traits to encourage or discourage certain uses may be an avenue for mitigating risk of negative effects for at-risk individuals (such as those with high neuroticism and signs of problematic use). Chatbots could potentially be tailored to different user types, with features and interaction styles that complement rather than replace human social interaction for vulnerable users.
For individuals with limited access to human social support, companion chatbots could serve as an intervention to reduce loneliness and its associated health risks. However, the design and implementation of these interventions must be carefully considered. Future HCI research should consider multiple pathways of influence when studying and designing AI companion technologies.
Ethical Considerations & Long-term Societal Impacts. The potential for some users to engage in deep, emotional relationships with chatbots raises ethical concerns. Developers and policymakers need to consider how to balance the potential benefits of AI companionship with the risks of over-reliance on artificial relationships. This may involve implementing safeguards or guidance systems within chatbots to encourage healthy usage patterns and human social interaction.
As AI technology advances and companion chatbots become more sophisticated, it is crucial to consider the long-term societal impacts of widespread AI companionship. Questions arise about how these technologies might shape social norms, expectations for human relationships, and the development of social skills, particularly for younger generations. Should we as a society push for less artificial companionship, or should we encourage a healthy mix of artificial and real companionship? Consider the case of a profoundly lonely and emotionally vulnerable individual. They might cling to a relationship that can provide any crumb of affection, no matter how unhealthy it becomes. Could companion chatbots fulfill that desperate need in a beneficial way, even if “unhealthy attachment” to the chatbot is part of the process? Perhaps alleviating loneliness enough to reduce the individual’s vulnerability could allow them to become more conscientious about seeking out and maintaining healthy, respectful human relationships.
5.4 Future Research Directions
Several key areas warrant further investigation. Longitudinal studies would be particularly valuable in understanding the causal relationships between chatbot usage, loneliness, and social outcomes over time. Such studies could track changes in social behaviors, relationship patterns, and psychological well-being as individuals engage with companion chatbots over extended periods.
Experimental studies manipulating specific aspects of chatbot design or usage could help identify features that promote healthy engagement while minimizing risks of problematic use or social displacement. These studies might explore factors such as conversation topics, interaction frequency, or anthropomorphic qualities of chatbots and their impacts on user experiences and outcomes.
Given the complex interactions observed in our model, future research should also investigate potential non-linear relationships between chatbot usage and psychological outcomes. This could involve more sophisticated statistical modeling techniques or qualitative approaches to capture nuanced patterns of usage and their effects.
The cultural context in which companion chatbots are used likely plays a significant role in shaping their impact. Cross-cultural studies could explore how different societal norms, communication styles, and attitudes towards technology influence the relationship between chatbot usage and loneliness.
Focused studies on specific user groups, such as the clusters we identified or specific demographic groups such as adolescents, older adults, or individuals with social anxiety, could yield valuable insights for developing targeted designs or guidelines.
As companion chatbots become more prevalent, it will be crucial to investigate their long-term effects on social skills, relationship expectations, and overall social functioning.
5.4.1 Limitations and Considerations
Survey design and recruitment. As our survey consists entirely of self-reported measures for usage characteristics and psychosocial measures, we face a considerable risk of recall bias. Our custom instruments for usage characteristics offered fairly broad options to reduce potential inaccuracies, though this also makes them coarse measures. Conducting a study involving monitoring or reporting of actual use would provide more specific and accurate usage data, and is in our plans for future work.
While our adaptations of established scales were based on factor loadings from previous studies and tailored to our research context, we acknowledge that these modifications may affect the scales’ validated psychometric properties. Future work should consider validating these adapted scales specifically for chatbot-related research. Due to the extensive nature of our survey, we prioritized reducing participant fatigue over maintaining the full length of each scale. This trade-off allowed us to explore a broader range of factors but may have impacted the depth of measurement for individual constructs. Additionally, our participants were primarily white/Caucasian individuals from the United States; applying our findings to other demographics should be done cautiously.
Clustering. We note that our clusters have their own set of limitations. While the number of clusters has statistical grounding, the final number can be considered somewhat arbitrary and based on our subjective judgments of interpretability. The heterogeneity of covariance matrices indicated by Box’s M-test requires caution in interpreting these results. Even so, this 7-cluster solution provides a starting point for further work on styles of use.
Regression Model. The exploratory nature of our model selection process warrants cautious interpretations and further investigation into the relationships it implies. The assumption of linear relationships may not capture complex nonlinear relationships in the population of users of companion chatbots. Further, due to the numerous predictors involved in the exploration and the final model, the risk of Type II errors is considerable [Lavery et al.(2019)]; our analysis may not capture all potential relationships of these variables.
Additionally, while we can theorize about the directions of relationships between variables, our cross-sectional design limits causal inferences. Future research employing experimental designs or longitudinal studies could provide more definitive evidence of causal relationships and explore potential non-linear effects. We plan to follow up with our participants to see changes in psychological, social, and usage characteristics at different time points.
@@ 6 Conclusion
This study examines the relationship between companion chatbot usage and loneliness, developing a model that explains approximately 50% of variance in loneliness through the complex interplay of psychological, social, and technological factors. Our findings reveal that while chatbot session length does not directly predict loneliness, this relationship is mediated by problematic use and moderated by social attraction to close others, highlighting the nuanced ways in which AI companions influence social well-being.
The model we developed through regression analysis identifies key mechanisms, including problematic use as a mediator and social attraction as a moderator, while accounting for personality traits such as neuroticism and social network characteristics. This suggests that the impact of companion chatbot usage on loneliness operates through complex pathways rather than direct effects.
Our cluster analysis further underscores this complexity, identifying seven distinct user profiles that demonstrate how similar usage patterns can lead to markedly different outcomes. For instance, socially fulfilled dependent users maintain healthy relationships despite intensive chatbot use, while lonely moderate users with similar usage patterns show signs of social withdrawal. These contrasting outcomes emphasize the importance of individual characteristics and social contexts in shaping the effects of AI companionship.
As AI companions become increasingly prevalent, these findings have crucial implications for design and deployment. The identification of distinct user profiles and key mediating factors suggests the need for personalized approaches and built-in mechanisms to detect and intervene in potentially harmful usage patterns. The contrasting outcomes between different user types raise important questions about how to design AI companions that support psychosocial well-being.
Future research should explore the causal relationships and long-term impacts of AI companionship through longitudinal studies while investigating how different design approaches might better support healthy engagement patterns. As we continue to navigate the evolving landscape of human-AI interaction in an increasingly digital world, understanding these complex relationships becomes crucial for developing technologies that genuinely enhance human well-being.
@@ 7 Acknowledgments
Statistical support was provided by data science specialist Jinjie Liu, at the Institute for Quantitative Social Science, Harvard University.
Appendix A Significant Correlations with Key Variables
Below, we include a list of significant (corrected $p<0.05$) correlations with chatbot usage frequency score, chatbot session length score, and loneliness (ULS-8). Negative correlations have text colored in blue. Additionally, the correlation matrix with original variable names is seen in Figure 18.
@@ Appendix A Significant Correlations with Key Variables
  • Conscientiousness (BFI-10): $r=-0.3351,p<0.001$
  • Socialization (LSNS): $r=-0.5055,p<0.001$
  • Perceived Social Support (MSPSS): $r=-0.6790,p<0.001$
  • Self-Esteem (B-RSES): $r=-0.6380,p<0.001$
@@
  • Perceived Chatbot Predictability: $r=-0.2981,p<0.001$
@@
  • Trust in Unfamiliar People (HHI): $r=-0.2045,p<0.001$
  • Willingness to Engage with Unfamiliar People (HHI): $r=-0.2536,p<0.001$
  • Trust in Familiar People (HHI): $r=-0.5132,p<0.001$
  • Willingness to Engage with Familiar People (HHI): $r=-0.5495,p<0.001$
  • State Empathy towards Close Person: $r=-0.3821,p<0.001$
  • Affective State Empathy towards Close Person: $r=-0.3725,p<0.001$
  • Cognitive State Empathy towards Close Person: $r=-0.3377,p<0.001$
  • Associative State Empathy towards Close Person: $r=-0.3382,p<0.001$
  • Social Attraction towards towards Close Person: $r=-0.2047,p<0.001$
  • Task Attraction towards towards Close Person: $r=-0.2755,p<0.001$
  • Perceived Similarity with Close Person: $r=-0.2670,p<0.001$
Appendix B Cluster Profiles
A heatmap of cluster means for variables at their original scales in each cluster, colored based on their deviation from the mean, can be seen in Figure 19. Figure 20 is a demographic table with percentages of each category for each cluster. Detailed descriptions for each cluster can be found below.
@@ Appendix B Cluster Profiles
Figure 20: Demographic table for study participants, including overall percentage for each category and each cluster.
B.0.1 Cluster 0: Disengaged Light Users
Disengaged light users (7.67% of sample) show low chatbot usage frequency ($z=-2.21$) and session length ($z=-1.09$), with average levels of loneliness ($z=-0.16$). This cluster represents users with minimal engagement in chatbot interactions.
This group demonstrates average extraversion ($z=0.80$) and neuroticism ($z=-0.37$), but low openness ($z=-1.94$). They exhibit average levels of perceived social support ($z=-0.01$) and self-esteem ($z=0.05$). Their trust towards chatbots ($z=-0.35$) and expectations from chatbot interactions ($z=-0.08$) are average, indicating a neutral view of AI technology.
@@ B.0.1 Cluster 0: Disengaged Light Users
B.0.2 Cluster 1: Well-Adjusted Moderate Users
Well-adjusted moderate users (23.02% of sample) show average chatbot usage frequency ($z=0.41$) and session length ($z=0.67$), with low levels of loneliness ($z=-1.22$). This cluster represents users who are socially fulfilled and use chatbots moderately.
This group demonstrates high extraversion ($z=1.08$), low neuroticism ($z=-1.24$), and high conscientiousness ($z=1.21$). They exhibit high levels of socialization ($z=1.46$), perceived social support ($z=1.26$), and self-esteem ($z=1.46$). Their trust towards chatbots ($z=0.93$) and expectations from chatbot interactions ($z=0.82$) are above average, indicating a positive view of AI technology. In general, they have above-average trust in people as well as attraction and empathy towards a close person. They tend to feel that interacting with chatbots makes talking to people somewhat easier and more enjoyable.
@@ B.0.2 Cluster 1: Well-Adjusted Moderate Users
B.0.3 Cluster 2: AI-Wary Light Users
AI-wary light users (11.88% of sample) show average chatbot usage frequency ($z=0.05$) and session length ($z=-0.45$), with low levels of loneliness ($z=-1.07$). This cluster may represent users who are cautious about AI technology.
This group demonstrates above-average openness ($z=1.33$), though their trust towards chatbots ($z=-2.02$), expectations from chatbot interactions ($z=-2.12$), and overall empathy and attraction towards a companion chatbot are very low, indicating skepticism towards AI technology and emotional distance from companion chatbots. They have above-average empathy and attraction towards a close person and feel that interacting with companion chatbots have little effect on their real-life relationships.
@@ B.0.3 Cluster 2: AI-Wary Light Users
B.0.4 Cluster 3: Lonely Moderate Users
Lonely moderate users (13.61% of sample) show average chatbot usage frequency ($z=0.44$) and session length ($z=-0.35$), with high levels of loneliness ($z=1.03$). This cluster may represent users who turn to companion chatbots due to loneliness.
This group demonstrates low levels of socialization ($z=-1.24$), with above-average trust towards chatbots ($z=0.65$) and expectations from chatbot interactions ($z=0.65$). They have slightly above-average empathy and attraction to both a companion chatbot and a close person, though they have below-average trust in unfamiliar people and reduced interest, ease, and amount of interaction with other people as a result of using chatbots.

### v3

@@ 5.3.1 Theoretical Contributions
Our findings extend beyond simple social displacement theories [85], demonstrating that companion chatbots serve different functions based on user characteristics and usage patterns. This aligns with the Parasocial Contact Hypothesis [86], which suggests media characters can provide meaningful social experiences. The mediating role of problematic use suggests a pathway where increased chatbot engagement creates more opportunities for unhealthy usage patterns, which in turn exacerbate loneliness. This aligns with compensatory internet use theory [87], where individuals with social deficits may turn to technology to fulfill unmet social needs. However, when this compensatory use becomes problematic—loss of control, negative mood when unable to access—it may paradoxically worsen the loneliness it was meant to address. Extensive chatbot use may not be harmful if problematic patterns are avoided, as evidenced by our Fulfilled Dependent Users who maintain high engagement without increased loneliness. For those with strong existing social networks, companion chatbots may serve as beneficial supplements to human social interaction. However, socially isolated individuals risk problematic use [17, 18], and we must consider the risk of chatbots replacing human interactions and potentially exacerbating loneliness. The key finding is that similar usage patterns can lead to markedly different outcomes depending on user characteristics. We also demonstrate how manual thematic analysis can be effectively combined with LLM-based theme extraction to achieve both interpretive depth and comprehensive coverage in large-scale qualitative research, offering a scalable approach for mixed-methods studies.
5.4 Limitations and Future Research
The self-reported measures of our survey risk recall bias. Our custom instruments offered broad options to reduce inaccuracies but resulted in coarse measures. Our adapted scales were based on factor loadings from previous studies but may affect validated psychometric properties. To reduce participant fatigue, we prioritized shorter scales over maintaining full lengths. Our participants were primarily white/Caucasian Americans, limiting generalizability. Additionally, while statistically grounded, cluster numbers involve subjective judgments of interpretability. The heterogeneity of covariance matrices indicated by Box’s M-test requires caution in interpretation. Our exploratory regression model selection warrants cautious interpretations as well. Assumption of linear relationships may not capture complex nonlinear relationships. The numerous predictors increase the risk of Type II errors [88]. Our cross-sectional design limits causal inferences.
Our triangulated approach combining manual and automated thematic analysis strengthened theme validation, but LLM-based analysis may miss subtle contextual nuances and interpretive depth that human analysts capture. The automated approach also relies on the quality and consistency of the underlying language model, which may introduce systematic biases in theme identification or prevalence estimation. Another limitation is that our study identifies vulnerable populations, but we did not assess whether participants were currently receiving mental health treatment or had been advised against using AI companions, which could affect the interpretation of our findings regarding vulnerable users.
Several key areas warrant investigation: longitudinal studies to understand causal relationships; experimental studies manipulating specific chatbot design aspects; investigation of non-linear relationships; cross-cultural studies; and research on long-term effects on social skills and relationship expectations.
@@ 6 Conclusion
This study examines the relationship between companion chatbot usage and loneliness, revealing complex interactions among psychological, social, and technological factors. Our model explaining approximately 50% of variance in loneliness demonstrates that while chatbot session length does not directly predict loneliness, this relationship is mediated by problematic use and moderated by social attraction to close others, with personality traits and social network characteristics serving as key confounding variables.
Our cluster analysis identified seven distinct user profiles demonstrating how similar usage patterns can lead to markedly different outcomes. Fulfilled Dependent Users maintain healthy relationships despite intensive chatbot use, while Lonely Moderate Users with similar patterns show signs of social withdrawal, emphasizing the importance of individual characteristics and social contexts in determining outcomes. Most participants used companion chatbots primarily for technological exploration and entertainment rather than social purposes, suggesting these tools serve broader functions than typically assumed. However, distinct vulnerable populations emerged who rely on chatbots for companionship and emotional support, raising important ethical considerations.
The ethical implications of our findings extend beyond individual user experiences to broader questions about technology’s role in human social development. The identification of distinct user profiles and key mediating factors suggests the need for personalized approaches and built-in mechanisms to detect potentially harmful usage patterns. As AI companions become increasingly sophisticated and widespread, the need for responsible development practices becomes more urgent. Our research provides an empirical foundation for moving beyond abstract ethical debates toward concrete, evidence-based approaches to AI companion design that prioritize user well-being and social health.
@@ 7 Acknowledgments
Statistical support was provided by data science specialist Jinjie Liu, at IQSS, Harvard University.
Appendix A Appendix
Appendix B Automated Thematic Analysis Implementation Details
Our automated thematic analysis employed OpenAI’s GPT-4 model with the following specifications:
  • Model: gpt-4o
  • Temperature: 0.2-0.3
  • Max tokens: 2000-3000 per response
The model was instructed to identify 5-10 key themes and provide: concise theme names, brief descriptions, representative quote IDs (2-3 per theme), and prevalence estimates (categorical: very common, common, moderate, rare).
The following prompt template was used for each question:
1You are analyzing anonymous survey responses to the question:
2"[QUESTION TEXT]"
3
4Here is a list of [N] responses:
5[NUMBERED RESPONSE LIST]
6
7Instructions:
81. Identify 5-10 key themes that appear across multiple responses.
92. For each theme, include:
10 - A concise theme name
11 - A short description
12 - 2-3 representative quotes (use quote numbers only)
13 - An estimate of prevalence (e.g., very common’, moderate’, rare’)
143. Do not include any identifying information.
15
16Format your output as JSON:
17
18{
19 "themes": [
20 {{
21 "theme_name": "Concise name",
22 "description": "Short description",
23 "quote_ids": [3, 9, 12],
24 "prevalence": "moderate"
25 }}
26 ]
27}
To examine how usage motivations and conversation topics varied across user clusters, we conducted a secondary analysis comparing theme prevalence within each cluster. Using the same GPT-4 pipeline, we analyzed free-text responses grouped by cluster membership, providing context about user characteristics and previously identified themes. For each of the six survey questions, responses were grouped by cluster assignment and analyzed separately. The LLM was provided with cluster-specific response sets, previously identified themes, and instructions to identify theme prevalence differences across clusters.
The following prompt template was used for each question:
1You are analyzing free-text responses to the question:
2
3"[QUESTION TEXT]"
4
5Here is a list of known themes identified in prior analysis:
6[THEME LIST WITH DESCRIPTIONS]
7
8The responses have been grouped by cluster based on user similarities.
9
10Your task:
111. For each cluster, estimate which of the above themes are most or least common.
122. Comment on clear differences in theme prevalence across clusters.
133. Highlight any surprising or unique usage patterns.
144. If applicable, mention any new themes not listed.
15
16Please present your analysis in the following format for each cluster:
17
18Cluster [n]:
19- Common themes:
20- Uncommon themes:
21- Differences:
22- Highlights:
23
24Here are the responses grouped by cluster:
25
26[CLUSTER-GROUPED RESPONSES]
Results from cluster-based LLM analysis were integrated with manual thematic analysis findings to identify differential theme prevalence across user types. This approach revealed how similar themes (e.g., "companionship seeking") manifested differently across clusters with varying psychological and social characteristics. For example, Lonely Moderate Users (Cluster 3) showed higher prevalence of emotional disclosure themes, while Well-Adjusted Moderate Users (Cluster 1) more commonly used chatbots for practical utility and stress relief.
The cluster comparison analysis validated our quantitative clustering by demonstrating qualitatively distinct usage patterns and motivations across identified user types, strengthening the interpretation of our seven-cluster solution.
@@ 92. For each theme, include:
Appendix C Model Details
The full formula used for the final multiple regression model is: Loneliness (ULS-8) $\sim$ Session Length Score + Social Attraction towards Chatbot + Problematic Use (GPIUS2) + Socialization (LSNS) + Social Attraction towards Close Person + Neuroticism + Agreeableness + Sex + Age + Session Length Score $\times$ Social Attraction towards Close Person.
Appendix D Significant Correlations with Key Variables
Below, we include a list of significant (corrected $p<0.05$) correlations with chatbot usage frequency score, chatbot session length score, and loneliness (ULS-8). Negative correlations have text colored in blue. Additionally, the correlation matrix with original variable names is seen in Figure˜6.
@@ Appendix D Significant Correlations with Key Variables
  • Conscientiousness (BFI-10): r=0.3351,p<0.001r=-0.3351,p<0.001
  • Socialization (LSNS): r=0.5055,p<0.001r=-0.5055,p<0.001
  • Perceived Social Support (MSPSS): r=0.6790,p<0.001r=-0.6790,p<0.001
  • Self-Esteem (B-RSES): r=0.6380,p<0.001r=-0.6380,p<0.001
@@
  • Perceived Chatbot Predictability: r=0.2981,p<0.001r=-0.2981,p<0.001
@@
  • Trust in Unfamiliar People (HHI): r=0.2045,p<0.001r=-0.2045,p<0.001
  • Willingness to Engage with Unfamiliar People (HHI): r=0.2536,p<0.001r=-0.2536,p<0.001
  • Trust in Familiar People (HHI): r=0.5132,p<0.001r=-0.5132,p<0.001
  • Willingness to Engage with Familiar People (HHI): r=0.5495,p<0.001r=-0.5495,p<0.001
  • State Empathy towards Close Person: r=0.3821,p<0.001r=-0.3821,p<0.001
  • Affective State Empathy towards Close Person: r=0.3725,p<0.001r=-0.3725,p<0.001
  • Cognitive State Empathy towards Close Person: r=0.3377,p<0.001r=-0.3377,p<0.001
  • Associative State Empathy towards Close Person: r=0.3382,p<0.001r=-0.3382,p<0.001
  • Social Attraction towards towards Close Person: r=0.2047,p<0.001r=-0.2047,p<0.001
  • Task Attraction towards towards Close Person: r=0.2755,p<0.001r=-0.2755,p<0.001
  • Perceived Similarity with Close Person: r=0.2670,p<0.001r=-0.2670,p<0.001
Appendix E Cluster Profiles
Figure 9: Significant pairwise differences between clusters on key variables based on Dunn’s test with Bonferroni correction. Cell background color indicates significance of differences (darker green = presence of greater significance in differences), with text color indicating direction (red = greater than, blue = less than). For example, >2>2*** indicates significantly greater values than Cluster 2 at $p<.001$. Significance levels: *: $p<.05$, **: $p<.01$, ***: $p<.001$.
Figure 10: Manual themes derived from stratified random sample (N=105) with corresponding LLM-detected themes from full dataset (N=404). Prevalence estimates in parentheses indicate LLM-based findings where they differ from manual assessment. Cluster variations show differential theme prominence across user types, with higher/lower prevalence clusters identified.
A heatmap of cluster means for standardized variables can be seen in Figure˜12, and one for variables at their original scales can be seen in Figure˜13. Figure˜11 is a demographic table with percentages of each category for each cluster. Detailed descriptions for each cluster can be found below.
@@ Appendix E Cluster Profiles
E.0.1 Cluster 0: Disengaged Light Users
Disengaged light users (7.67% of sample) show low chatbot usage frequency (z=2.21z=-2.21) and session length (z=1.09z=-1.09), with average levels of loneliness (z=0.16z=-0.16). This cluster represents users with minimal engagement in chatbot interactions.
This group demonstrates average extraversion ($z=0.80$) and neuroticism (z=0.37z=-0.37), but low openness (z=1.94z=-1.94). They exhibit average levels of perceived social support (z=0.01z=-0.01) and self-esteem ($z=0.05$). Their trust towards chatbots (z=0.35z=-0.35) and expectations from chatbot interactions (z=0.08z=-0.08) are average, indicating a neutral view of AI technology.
@@ E.0.1 Cluster 0: Disengaged Light Users
E.0.2 Cluster 1: Well-Adjusted Moderate Users
Well-adjusted moderate users (23.02% of sample) show average chatbot usage frequency ($z=0.41$) and session length ($z=0.67$), with low levels of loneliness (z=1.22z=-1.22). This cluster represents users who are socially fulfilled and use chatbots moderately.
This group demonstrates high extraversion ($z=1.08$), low neuroticism (z=1.24z=-1.24), and high conscientiousness ($z=1.21$). They exhibit high levels of socialization ($z=1.46$), perceived social support ($z=1.26$), and self-esteem ($z=1.46$). Their trust towards chatbots ($z=0.93$) and expectations from chatbot interactions ($z=0.82$) are above average, indicating a positive view of AI technology. In general, they have above-average trust in people as well as attraction and empathy towards a close person. They tend to feel that interacting with chatbots makes talking to people somewhat easier and more enjoyable.
@@ E.0.2 Cluster 1: Well-Adjusted Moderate Users
E.0.3 Cluster 2: AI-Wary Light Users
AI-wary light users (11.88% of sample) show average chatbot usage frequency ($z=0.05$) and session length (z=0.45z=-0.45), with low levels of loneliness (z=1.07z=-1.07). This cluster may represent users who are cautious about AI technology.
This group demonstrates above-average openness ($z=1.33$), though their trust towards chatbots (z=2.02z=-2.02), expectations from chatbot interactions (z=2.12z=-2.12), and overall empathy and attraction towards a companion chatbot are very low, indicating skepticism towards AI technology and emotional distance from companion chatbots. They have above-average empathy and attraction towards a close person and feel that interacting with companion chatbots have little effect on their real-life relationships.
@@ E.0.3 Cluster 2: AI-Wary Light Users
E.0.4 Cluster 3: Lonely Moderate Users
Lonely moderate users (13.61% of sample) show average chatbot usage frequency ($z=0.44$) and session length (z=0.35z=-0.35), with high levels of loneliness ($z=1.03$). This cluster may represent users who turn to companion chatbots due to loneliness.
This group demonstrates low levels of socialization (z=1.24z=-1.24), with above-average trust towards chatbots ($z=0.65$) and expectations from chatbot interactions ($z=0.65$). They have slightly above-average empathy and attraction to both a companion chatbot and a close person, though they have below-average trust in unfamiliar people and reduced interest, ease, and amount of interaction with other people as a result of using chatbots.